Circular No. 7215

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999da IN NGC 6411 R. Johnson and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery of an apparent supernova with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1999da was discovered on unfiltered images taken on July 5.4 UT (mag ∼ 17.0) and confirmed on images taken on July 2.4 (mag ∼ 17.6). The new object is located at α = 17^{h}35^{m}22^{s}.96, δ = +60^{o}48′49′′.3 (equinox 2000.0), which is 71′′.3 west and 1′′.2 north of the nucleus of NGC 6411. A KAIT image of the same field on June 29.4 (limiting mag ∼ 19.5) showed nothing at the position of SN 1999da.

SUPERNOVA 1998bw IN ESO 184-G82 F. Patat, European Southern Observatory (ESO); E. Cappellaro, L.

Rizzi, and M. Turatto, Padua Observatory; and S. Benetti, Telescopio Nazionale Galileo, report on behalf of the ESO Team (cf. IAUC 6895): “We have obtained broadband BVRI photometry of SN 1998bw using the ESO/Dutch 0.9-m telescope and the ESO 3.6-m telescope (+ EFOSC2) at La Silla on four dates between Mar. 16 and June 17 (i.e., between 310 and 403 days after B maximum). The lightcurve does not appear to deviate significantly from the exponential decay already discussed by McKenzie and

Schaefer (http://xxx.lanl.gov/abs/astro-ph/?9904397) for the phase

range 50–270 days. The decline rates in magnitudes per 100 days, computed using our new data and the data by Galama et al. (1998, Nature 395, 670) and McKenzie and Schaefer are 1.52 ± 0.01 in B, 1.73 ± 0.01 in V, 1.63 ± 0.04 in R, and 1.64 ± 0.03 in I. With the exception of the B band, all decline rates appear to have slightly decreased with respect to the values reported by McKenzie and Schaefer. Only for the data taken on June 16 does there seem to be a lightcurve flattening, especially in R and I, which could be the first indication that the ejecta/circumstellar-matter interaction has commenced. This seems to be confirmed by a spectrum that we took at the ESO 3.6-m reflector (+ EFOSC2) on May 21.2 UT (range 335–1000 nm, resolution 1.7 nm). A preliminary analysis of the data shows that the Mg I] line at 457.1 nm has an asymmetric profile and at least two components, with FWHM ∼ 8000 and 2000 km/s. Similar results are obtained for the [O I] doublet at 630.0 and 636.4 nm. The narrow component could be generated in a fast expanding wind shocked by the supernova ejecta. The overall morphology of the spectrum is very similar to that of a type-Ib supernova at a similar phase.”

1999 July 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green